Mix 12 units of oxytocin in 500 mL NS. Start infusion at 18 mL/hr. Increase by 5 mL/hr every 10 minutes until uterine contractions are 2 to 3 minutes apart. Do not exceed 32 milliunits/min. Maintain total IV fluid volume at 100 mL/hr. How many milliunits per minute will be delivered when you start this infusion?(round to the nearest tenth)
Answer #1

This infusion is initially started at 18 ml/hour.

So per minute = 18/60 = 0.3 ml/minute

12 units present in 500 ml.

So in 500 ml, 12000 miliunits are present.

In 0.3 ml , 12000/5000 * 0.3 = 0.72 miliunits.

So 0.72 miliunits will be delivered per minute when the infusion is started.